1)Find the limit.
a:)lim t->0 (e^(-3t)i+(t^2/sin^2 t)j|+ cos 2tk)
b:)lim t->1((t^2-t)/(t-1)i+sqrt(t+ 8)j+(sin pi t)/(ln t) k)

2) . Find a vector equation and parametric equations for the line segment that joins P(-1,2,2) and Q(-3,5,1).
1
Expert's answer
2020-09-15T17:28:35-0400

1)

a) "lim_{ t \to 0} (e^{-3t}i+\frac{t^2}{sin^2 t}j+ cos 2tk)=(lim_{ t \to 0} e^{-3t})i+(lim_{ t \to 0}\frac{t^2}{sin^2 t})j+ (lim_{ t \to 0} cos 2t)k."

Now we can find each limit:

"lim_{ t \to 0} (e^{-3t})=e^{-3\cdot 0}=e^0=1".

To find the second limit we can use L'Hopital's rule twice:

"lim_{ t \to 0}(\frac{t^2}{sin^2 t})=[\frac{0}{0}]=lim_{ t \to 0}(\frac{(t^2)\prime}{(sin^2 t)\prime})=lim_{ t \to 0}(\frac{2t}{2sint cos t})=lim_{ t \to 0}(\frac{2t}{sin2t})=[\frac{0}{0}]=1".

"lim_{ t \to 0} (cos 2t)=cos(2\cdot 0)=cos0=1."

Finally:

"lim_{ t \to 0} (e^{-3t}i+\frac{t^2}{sin^2 t}j+ cos 2tk)=1i+1j+1k=i+j+k."


b) "lim_{t \to 1}(\frac{t^2-t}{t-1}i+\sqrt{t+ 8}j+\frac{sin \pi t}{ln t} k)=(lim_{t \to 1}\frac{t^2-t}{t-1})i+(lim_{t \to 1}\sqrt{t+ 8})j+(lim_{t \to 1}\frac{sin \pi t}{ln t}) k."

Now we can find each limit:

"lim_{t \to 1}(\frac{t^2-t}{t-1})=[\frac{0}{0}]=lim_{t \to 1}(\frac{t(t-1)}{t-1})=lim_{t \to 1}t=1."

"lim_{t \to 1}\sqrt{t+ 8}=\sqrt{1+ 8}=\sqrt{9}=3."

"lim_{t \to 1}\frac{sin \pi t}{ln t}=[\frac{0}{0}]=lim_{t \to 1}\frac{(sin \pi t)\prime}{(ln t)\prime}=lim_{t \to 1}\frac{\pi cos \pi t}{\frac{1}{t}}=lim_{t \to 1}(\pi t cos \pi t)="

"=\pi \cdot 1 \cdot cos(\pi \cdot 1)=- \pi."

Finally:

"lim_{t \to 1}(\frac{t^2-t}{t-1}i+\sqrt{t+ 8}j+\frac{sin \pi t}{ln t} k)=1i+3j+\pi k=i+3j-\pi k."


2)

The vector equation of the line segment is given by "r(t)=(1-t)r_0+tr_1," where "0\leq t\leq 1" and "r_0" and "r_1" are the vector equivalents of the endpoints:

"P(-1,2,2)" becomes "r_0=\u27e8-1,2,2\u27e9,"

"Q(-3,5,1)" becomes "r_1=\u27e8-3,5,1\u27e9."

Plug "r_0" and "r_2" into the vector formula:

"r(t)=(1-t)\u27e8-1,2,2\u27e9+t\u27e8-3,5,1\u27e9=\u27e8-1+t-3t,2-2t+5t,2-2t+t\u27e9="

"=\u27e8-1-2t,2+3t,2-t\u27e9".

We have the vector equation: "r(t)=⟨-1-2t,2+3t,2-t⟩, 0 \leq t \leq 1," or "r(t)=(-1-2t)i+(2+3t)j+(2-t)k, 0 \leq t \leq 1."

To get the parametric equations of the line segment, we have to take the coefficients by "i, j, k":

"\begin{cases} x=-1-2t, \\ y=2+3t, \\ z=2-t, \end{cases}" "0 \leq t \leq 1."